OpenStudy (anonymous):

Andrea can clean a house 4 times as fast as Denise. When they work together, Andrea and Denise can clean a large house in 8 hours. How many hours would it take Denise to clean the house by herself? 20 15 10 5

OpenStudy (anonymous):

@lalaly

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

ANDREA'S one day work =1/4

OpenStudy (anonymous):

ok but that is not one the answer choices! @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

it is a bit confusing.

OpenStudy (cggurumanjunath):

@jdoe0001

OpenStudy (anonymous):

these are the answers 20 15 10 5

OpenStudy (lalaly):

Sorry i found my mistake give me a min

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

so let's see how much they did in 1hr each \(\large \begin{array}{cccccc} denise&&andrea&&total\\ \hline\\ \cfrac{1}{d}&+&\cfrac{1}{{\color{blue}{ a}}}&=&\cfrac{1}{8}&\implies \cfrac{1}{d}+\cfrac{1}{{\color{blue}{ 4d}}}=\cfrac{1}{8} \end{array}\)

OpenStudy (anonymous):

@lalaly did you find the mistake!

OpenStudy (anonymous):

ok then what should I do next? @jdoe0001

OpenStudy (jdoe0001):

well, solve for "d", which is how long "denise" takes

OpenStudy (jdoe0001):

the idea being that we dunno how long they each really take in 1hr andrea can do in 1hr 1/(nth) of the work and that (nth) figure, will depend on her speed, so her work rate is 1/(nth) => 1/a same goes for "denise", she'll do 1/(nth) of the work in 1hr, thus 1/d since "a" and "d" are two distinct values because andrea is faster than denise but we do know that since they finish the whole house in 8hrs they can only do 1/8th of the work in 1hr

OpenStudy (anonymous):

ok and then @jdoe0001

OpenStudy (jdoe0001):

and then you'd solve for "d"

OpenStudy (jdoe0001):

since denise's rate is 1/d in 1hr "d" is how long she really takes by herself

OpenStudy (anonymous):

so the answer would be B @jdoe0001

OpenStudy (jdoe0001):

well, what did you get?

OpenStudy (anonymous):

Wait for the 4d what do I multiply the 4 with?? @jdoe0001

OpenStudy (jdoe0001):

you may want to multiply both sides by "8d" see what that gives you

OpenStudy (anonymous):

what i got was 8d/8d + 8d/32d = 1/64d @jdoe0001

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ 8d\times }}\cfrac{1}{d}+{\color{blue}{ 8d\times }}\cfrac{1}{4d}={\color{blue}{ 8d\times }}\cfrac{1}{8}\implies \cfrac{8d}{d}+\cfrac{8d}{4d}=\cfrac{8d}{8}\) can we cancel out anything terms there from the fractions?

OpenStudy (anonymous):

the d! @jdoe0001

OpenStudy (jdoe0001):

\(\bf \cfrac{8\cancel{d}}{\cancel{d}}+\cfrac{\cancel{8d}}{\cancel{4d}}=\cfrac{\cancel{8}d}{\cancel{8}}\)

OpenStudy (jdoe0001):

so.. what do you think we're left with?

OpenStudy (anonymous):

8 =d

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... \(\bf \cfrac{8d}{4d}\implies \cfrac{8}{4}\cdot \cfrac{d}{d}\) so... what would our \(\cfrac{8}{4}\) turn into?

OpenStudy (jdoe0001):

the 1st fraction gives 8, yes the 2nd gives something else

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

ahhhh so we end up with \(\bf {\color{blue}{ 8d\times }}\cfrac{1}{d}+{\color{blue}{ 8d\times }}\cfrac{1}{4d}={\color{blue}{ 8d\times }}\cfrac{1}{8}\implies \cfrac{8d}{d}+\cfrac{8d}{4d}=\cfrac{8d}{8}\implies 8+2=d\) now you can see how long "denise" takes by herself :)

OpenStudy (anonymous):

ITS 10! OH OK! @jdoe0001

OpenStudy (jdoe0001):

hmmm lemme rethink that one for a sec...

OpenStudy (jdoe0001):

somehow it has a bad taste =)

OpenStudy (anonymous):

oh ok! @jdoe0001 can you help me on another one!

OpenStudy (jdoe0001):

hmmm something tells me that 10 is Andrea's =(

OpenStudy (jdoe0001):

hehhe

OpenStudy (jdoe0001):

let's see andrea is 4 times as fast as denise so whatever denise speed is, andrea's has to be 4 times that much

OpenStudy (jdoe0001):

lemme quickly rewrite it

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

are those the only choices you have? I seem to be getting 40

OpenStudy (anonymous):

Yeah those are the only ones!

OpenStudy (jdoe0001):

hmmm you see, it can't be 10hrs shoot if both take 8hrs together, and andrea is faster andrea would take about 10hrs denise would take 4 times 10, or 40

OpenStudy (jdoe0001):

each one will take longer than 8hrs because would be working alone

OpenStudy (jdoe0001):

the only way I can see one of the choices, is if andrea is only 2 times as fast

OpenStudy (jdoe0001):

so.... it can't be 10 for denise because that'd meab 1/4 of 10 for andrea meaning 2.5hrs so andrea would finish the whole house in 2.5hrs while working with denise, will take 8hrs lol heheh so that'd imply andrea is a speeding bullet by herself, but with denise she's bogged down =)

OpenStudy (jdoe0001):

but I'd say based on the info provided, andrea being 4 times as fast I'd only get 40hrs for denise

OpenStudy (anonymous):

@jdoe0001 last one please Two bricklayers, Sam and Joe, are working on your house. Sam can complete the work in 5 hours, while Joe can complete the work in 3 hours. How many hours does the bricklaying take if they work together? 15/8 15/9 6/8 4/3

OpenStudy (anonymous):

@mary.rojas

OpenStudy (anonymous):

the one all the way at the bottom! @mary.rojas

OpenStudy (anonymous):

the one that says about two bricklayers @jdoe0001

OpenStudy (jdoe0001):

same layout pretty much, Sam in 1 hr can do 1/5 of the work Joe 1hr can do 1/3 of the work that's their work rate in 1 hr they'll do 1/t of the work if they put their RATE together, how long will it take? well \(\large \large \begin{array}{cccccc} sam&&joe&&total\\ \hline\\ \cfrac{1}{5}&+&\cfrac{1}{3}&=&\cfrac{1}{t} \end{array}\) they'll finish together at "t" time

OpenStudy (jdoe0001):

now if we multiply both sides by "15t" then we end up with \(\bf 15t\times \cfrac{1}{5}+15t\times \cfrac{1}{3}=15t\times \cfrac{1}{t}\implies \cfrac{15t}{5}+\cfrac{15t}{3}=\cfrac{15t}{t} \\ \quad \\ 3t+5t=15\)

OpenStudy (anonymous):

5t=15 t=3

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, 3t+5t = 8t so 3t+5t=15 => 8t = 15 then divide both sides by 8 :)

OpenStudy (anonymous):

15/8 @jdoe0001

OpenStudy (jdoe0001):

15/8 \(\checkmark\)

OpenStudy (anonymous):

@jdoe0001 What is the simplified form of 4x^2-25/2x-5 2x - 5, with the restriction x≠-5/2 2x + 5, with the restriction x≠5/2 2x - 5, with the restriction x ≠ 5/ 2 2x + 5, with the restriction x ≠ -5/ 2

OpenStudy (anonymous):

I got 2x+5 but I am not sure of the restriction. I believe it is the second one!

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Is it B? @jdoe0001

OpenStudy (jdoe0001):

is a rational so restrictions will apply no to make it undefined meaning the denominator cannot be 0 meaning "x" shouldn't take anything that makes the denominator 0 so if you set the denominator to 0 and solve for "x", that'd give it out

OpenStudy (jdoe0001):

2x-5 = 0 => x = 5/2, thus \(\bf x\ne \frac{5}{2}\)

OpenStudy (whpalmer4):

Much simpler way of working the Andrea and Denise problem: Denise's rate is 1 house in \(d \text{ hours} = \dfrac{1}{d}\) Andrea's rate is 4x Denise's rate, or \(4*\dfrac{1}{d} = \dfrac{4}{d}\) Their combined rate is \(\large\frac{1}{d}+\frac{4}{d} = \frac{5}{d}\) 1 house in 8 hours working together means \[\frac{1}{\frac{5}{d}} = 8\]or\[\frac{d}{5}=8\]\[d=40\]So Denise takes 40 hours to do 1 house at her speedy pace of 1/40 house/hour. Andrea works 4 times faster, so she does 4 houses in 40 hours or 1 house in 10 hours, giving her a rate of 1/10 house/hour. I agree that none of the available answers are correct for the problem as stated.